OpenStudy (anonymous):

The Riemann sum, the limit as the maximum of delta x sub i goes to infinity of the summation from i equals 1 to n of f of the quantity x star sub i times delta x sub i , is equivalent to the limit as n goes to infinity of the summation from i equals 1 to n of f of the quantity a plus i times delta x, times delta x with delta x equals the quotient of the quantity b minus a and n. Write the integral that produces the same value as the limit as n goes to infinity of the summation from i equals 1 to n of the product of the quantity 2 plus 3 times i over n and 3 over n.

OpenStudy (anonymous):

https://gyazo.com/4a2702de2f727a0265591e6cfd66d3a9 @jim_thompson5910

OpenStudy (anonymous):

Same rules correct?

jimthompson5910 (jim_thompson5910):

do you see what the value of 'a' is?

OpenStudy (anonymous):

a=2 b=5

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now do you see what f(x) must be?

OpenStudy (anonymous):

It should be just x Final answer should be D

jimthompson5910 (jim_thompson5910):

I can't see your answer choices, but you'll find that a = 2 b = 5 f(x) = x so \[\Large \lim_{n \to \infty}\sum_{i=1}^{n}\left(2+\frac{3i}{n}\right)*\frac{3}{n} = \int_{2}^{5}x dx\]

OpenStudy (anonymous):

https://gyazo.com/f9c6a5061ff442c9b97d4c357dd71289 :)

jimthompson5910 (jim_thompson5910):

yep it's D

OpenStudy (anonymous):

Is it cool if I just post two other questions here so you can just check them? Instead of making other questions?

jimthompson5910 (jim_thompson5910):

sure

jimthompson5910 (jim_thompson5910):

just as long as it doesn't get too cluttered

OpenStudy (anonymous):

If either one is wrong, I'll make a new question.

jimthompson5910 (jim_thompson5910):

They're both correct. Nice job

OpenStudy (anonymous):

Appreciate it man

jimthompson5910 (jim_thompson5910):

no problem